OpenStudy (anonymous):

The gum you like to buy is on sale. It is regularly priced at $1.59 Write an equation that will help you determine how much you will save if you buy the gum today. lets say s= sales price per pack c= cost savings

OpenStudy (deadly_roses):

1.59 - s = c

OpenStudy (idealist10):

Answer: 1.59-s=c

OpenStudy (deadly_roses):

If you want to know how much you will save, take the regular price and subtract the sale price. For example: If a CD normally costs $10 and today it is on sale for $8 you find your savings by doing: 10-8

OpenStudy (deadly_roses):

once you do that put your answer into the form i gave you: 1.59 - s = c

OpenStudy (anonymous):

Thank you I am really bad at math.